### v1

# MoSE: Skill-by-Skill Mixture-of-Expert Learning for Autonomous Driving

Abstract
Recent studies show large language models (LLMs) and vision language models (VLMs) trained using web-scale data can empower end-to-end autonomous driving systems for a better generalization and interpretation. Specifically, by dynamically routing inputs to specialized subsets of parameters, the Mixture-of-Experts (MoE) technique enables general LLMs or VLMs to achieve substantial performance improvements while maintaining computational efficiency. However, general MoE models usually demands extensive training data and complex optimization. In this work, inspired by the learning process of human drivers, we propose a skill-oriented MoE, called MoSE, which mimics human drivers’ learning process and reasoning process, skill-by-skill and step-by-step. We propose a skill-oriented routing mechanism that begins with defining and annotating specific skills, enabling experts to identify the necessary driving competencies for various scenarios and reasoning tasks, thereby facilitating skill-by-skill learning. Further align the driving process to multi-step planning in human reasoning and end-to-end driving models, we build a hierarchical skill dataset and pretrain the router to encourage the model to think step-by-step. Unlike multi-round dialogs, MoSE integrates valuable auxiliary tasks (e.g. description, reasoning, planning) in one single forward process without introducing any extra computational cost. With less than 3B sparsely activated parameters, our model outperforms several 8B+ parameters on CODA AD corner case reasoning task. Compared to existing methods based on open-source models and data, our approach achieves state-of-the-art performance with significantly reduced activated model size (at least by $62.5\%$) with a single-turn conversation.
1 Introduction
Figure 1: Activated model size and performance of various open-source specialist models on the CODA dataset. The proposed MoSE achieves SOTA performance with less than 3B activated parameters.
Figure 2: Overall model pipeline: (a) the step-by-step process of the model, where it progresses from perception to prediction and then to planning based on the input image and question. (b) the internal structure of each MoSE layer, where skill routing enables skill-by-skill learning. For each task, the model ultimately constructs a skill chain, which partially reflects its reasoning process, as shown in (c).
Systems with human-level understanding of multi-modal scenarios have become increasingly essential for highly intelligent applications. Many downstream tasks like autonomous driving and robotics, are craving Large Language Model (LLMs) or Vision Language Model (VLMs) to not only boost the overall performance, but also to improve the system’s interpretability, reasoning capabilities and interactive abilities. However, while web-scale pre-trained LLMs and VLMs excel at general knowledge, they often lack domain-specific knowledge required for particular tasks. Domain-specific models often require substantial design effort to effectively align task-specific details with the general knowledge encoded in pre-trained LLMs in order to achieve improved performance.
In the field of autonomous driving, some prior studies have presented promising advancements. For example, LMDrive [31] exhibits promising results on trajectory prediction task by training with instruction-trajectory data collected from the Carla simulator. The model focuses on low-level planning while still having large potential to integrate reasoning or interactive functionalities. OmniDrive [35] also focuses on reasoning task using related question-answering (QA) pairs. It covers questions for multiple subtasks, such as perception, prediction, and planning, integrating the trajectory prediction task as one of the QAs. The model can provide more interpretable decisions, but often suffers inconsistency issues between questions and faces gap between high-level and low-level planning results, leading to additional design efforts for alignment [17]. DriveLM [33] jointly performing Graph VQA and end-to-end driving. Based on the human-written reasoning logic between perception, prediction, planning, behavior and motion; the proposed GVQA dataset facilitate the full stack of driving tasks and serves as a comprehensive dataset for driving with language. CODA[5] adopts a hierarchical data structure and prompts powerful VLMs to analyze real world complex driving scenarios. The annotations from VLMs are then verified by human annotators. Different from DriveLM, which is built upon NuScenes[3], CODA primarily considers driving corner cases, especially object-level corner cases.
A key issue for domains that demand both high speed and cost efficiency, such as autonomous driving, is that large pre-trained models often struggle due to their substantial model size at inference. To address this, an emerging technique for pre-trained LLM/VLM is the sparse Mixture-of-Experts (MoE) layers in Transformer-based models [32]. They are designed to improve model efficiency and performance by dynamically activating different subsets of specialized sub-models, or “experts” for each input. Sparse MoE layers can substantially scaling up LLMs’ capacity with minimal computational overhead [20, 12, 18]. A router network learns to route each input to the most relevant experts, thereby enabling the model to handle a wide variety of tasks or data domains while using only a fraction of its total parameters for each input. MoE-LLaVA[23] demonstrates a scaling up performance of smaller-scale VLMs ($<$3B) using MoE technique to achieve competitive performance compared to 7B models on general visual language tasks. These advancements motivated us to explore feasible solutions for downstream tasks such as autonomous driving, where large-scale and diverse training data are essential but expensive to obtain. Consequently, a straightforward MoE approach encounters challenges and often results in suboptimal performance due to the limited data for specific downstream tasks.
In this paper, we explore the MoE technique in task-specific, small-scale ($<$3B), multi-modal vision-language models (VLMs) then introduce an efficient framework designed for autonomous driving tasks (c.f. Fig. 2). Inspired by human learning strategy, we design a skill-by-skill learning approach for the routing procedure with a skill centric routing strategy. The routing strategy identifies essential driving skills required for various driving scenarios and driving stages, guiding experts within each MoE layer to specialize in distinct scenarios and progressively develop the skills needed to address targeted conditions effectively.
Furthermore, inspired by progressive human driving behavior and prior end-to-end frameworks such as UniAD [15], we further employ a hierarchical routing strategy across layers to enable the system to think and operate in a step-by-step manner. This approach not only improves consistency in question-answering across varying levels but also aligns with valuable auxiliary tasks (e.g. perception, prediction and planning) in a single forward pass, eliminating the need for additional context or multi-round QA. Experimental results demonstrate that, with only 2B+parameters size (e.g. Qwen2VL-2B [34]), our model achieves performance competitive to other state-of-the-art autonomous driving reasoning tasks with more than 8B+parameters. To sum up, our contribution can be summarized as follows:
  • We propose a novel Mixture-of-Skill-Experts (MoSE) to enhance the reasoning capability of small-scale VLMs ($<$3B parameters) for autonomous driving. Inspired by the learning process of human drivers, our approach provides more interpretable reasoning capabilities.
  • We propose a skill-oriented routing mechanism, enabling the model to reason step-by-step and guiding experts learn scenario-specific driving skill-by-skill.
  • Compare to existing methods based on open-source models and data, MoSE achieves state-of-the-art performance on AD corner case reasoning task with significantly reduced activated parameters size (at least by $62.5\%$).
2 Related Work
LLM/VLM for Autonomous Driving
Some of the early explorations show their superior performance focusing on scene understanding and description tasks [28, 30, 36], which are build on close source models e.g. GPT3.5 and GPT4-v. Some work like LLM-Driver [6] leverage the LLM for AD task as a multi-task learner to predict both language and control signal. Due to the rapid development of multi-modal models since LLaVA [26, 25, 27], many works try different paths to leverage VLMs for AD building upon a large scale novelly annotated data [31, 33, 35]. LMDrive and OmniDrive use task tailored vision encoder instead of general ViT like [40], and align the extracted features [22] to LLMs. DriveLM proposed a Graph VQA dataset builds upon nuScences [3] and Carla simulator, and trains VLMs using driving related QA. The proposed agents take multi-view images as input and use different models for QA tasks and trajectory prediction tasks.
General Mixture-of- Experts
Mixture-of-experts architectures with sparse activations of experts provides a possible way to scale up LLMs with high efficient training and inference. In the transformer based architectures, one promising way is to substitute feed-forward network(FFN) layers with sparse MoE Layers [32, 43, 12]. The efficiency mainly comes from sparse activation mechanism, Top-1 [12] or Top-K [20] experts are activated by the routers at each MoE layer. Some works start to further enhance the training and inference efficiency through modified router. Lory proposed token batching mechanism to build an efficient and fully differentiable MoE. DeepSeekMoE [10] segments experts into finer groups and isolates certain experts as shared ones. This design allows for capturing common knowledge while enabling other experts to specialize, thereby improving performance and computational efficiency. Zhang et al. [42] proposed a method to diversify expert knowledge by grouping similar experts and pruning redundancies. This approach enhances parameter efficiency and ensures that each expert group captures distinct aspects of the data, leading to better model performance.
Multimodal Mixture-of-Experts
MoE-LLaVA [23] build lightweight VLMs using MoE and gain comparable performance on various visual understanding tasks with less half parameters. [7] automatically determines the number of activated experts for each task, avoiding the laborious manual tuning of optimal model size. [39] proposed to flexibly incorporate arbitrary modality combinations while maintaining robustness to missing data. The router specializes in handling fewer modality combinations by assigning the Top-1 gate to the expert corresponding to the observed modality combination. Med-MoE [19] introduced a lightweight framework for multi-modal medical task, and addressing both discriminative and generative needs in medical. They proposed to pre-train the rooter and apply domain-specific expert fine-tuning. Modality-aware MoE [24] divides experts into modality-specific groups (one group for text, one for vision) so that each group can focus on the features of its modality while still allowing cross-modal interactions in shared layers.
3 Method
In this section, we present the proposed Mixture-of-Skill-Experts method. Given a visual signal $I_{v}$ and a user instruction $I_{t}$, we develop a MoSE based VLM $F(\cdot)$ that generates driving reasoning result:
$R_{t}=F(I_{v},I_{t}).$ (1)
Input visual signal $I_{v}$ can be single-view or multi-view images or videos, while $I_{t}$ encompasses mostly questions about driving perception, prediction, planning, reasoning and suggestion.
As illustrated in Fig. 2(a), the overall system processes multimodal inputs, comprising visual observations in image format and language-based interactive inputs in text format, and generates a text-based output. We adopt general MoE into skill-by-skill learning at both the case and task levels, each task is decomposed into hierarchical levels, and for each case, different skills are activated at different levels (Sec. 3.1.1). We begin with language-visual-aligned, instruction-tuned, small-scale VLMs, inserting MoE layers, then initiating the router pre-training with a small amount of labeled data (Sec. 3.1.2 and  3.1.3). This step enables the router to focus on skill routing rather than token-level distinctions. The router learns to characterize necessary skills in each stage and activate appropriate skill experts. After the router training, we conduct supervised fine-tuning on the entire model (Sec. 3.2). Following the approach of Qwen2-VL, we employ a pre-trained Vision Transformer [11] (ViT) as the vision encoder, maintaining the existing processing pipeline for vision tokens. During training, we only finetune the MoSE blocks while keeping the rest of the model frozen.
3.1 Skill Centric Routing Mechanism
In a general MoE, a router network learns to select activated experts for each individual token automatically. The trained router tends to focus on distinguish different domains or modalities which may be inadequate for some tasks like autonomous driving. Compared to internet-scale data, autonomous driving images and videos exhibit higher similarity, with questions following specific patterns. These factors introduce additional challenges for MoE training, necessitating a carefully designed architecture and learning strategy to address limitations in data scale and diversity.
To enable skill-by-skill learning, the router must gain a refined understanding of both the driving scene and input text, facilitating expert selection for each necessary stage based on the overall driving context. To mimic human reasoning and learning process, we design a skill centric routing mechanism to guide the model learn and think skill-by-skill. To build such procedure, we start with define essential skills for our task, autonomous driving. One thing need to be mentioned is that the skills can be shared across different dataset. After having the skills, we perform data annotation using rule based method or other large VLMs, in our case, GPT-4o. The data quality is directly related to skill definition, so one can optimize the human annotation effort by adjusting the skills. After having the skill data, we pretrain the router of all target layer, the trained router can guide the expert follow a hierarchical path and generate answer step-by-step.
Figure 3: The process of constructing skill data consists of two stages. In the first stage, we define the skills relevant to the current task, specifically autonomous driving. In the second stage, we perform data annotation using both rule-based and VLM-based methods.
3.1.1 Skill Definition
To construct such a skill-centric routing mechanism, we start by defining driving skills. In doing so, we follow two key principles:
  1. 1.
    Considering the parameters introduced by the router, the complexity should be minimized to simplify routing.
  2. 2.
    The defined skills should cover all data while being applicable across multiple datasets.
To meet these, we introduce a general skill at each hierarchical level. On one hand, this general skill ensures comprehensive skill coverage, especially for relatively ambiguous samples. On the other hand, since some samples may not require all hierarchical levels, we assign general skills to the unused levels to ensure coverage for those samples.
The overall skill definition pipeline can be seen in Fig. 3 Stage 1, we leverage existing state-of-the-art large models e.g. GPT, to define the driving skills. During the definition process, we provide GPT with examples from different datasets leveraging its inherent world knowledge. We additionally incorporate human driving experience as supplementary information. Finally, GPT generates skill definitions at different hierarchical levels, and the skills are further refined by human. Taking the image and question in Fig. 2(a) as an example, we annotate the activated skills at each hierarchical level.
Our model attends to different hierarchical levels at different stages and activates the corresponding skills at each level. Under this skill definition, the activated skills across hierarchical levels form a structured chain, one possible structured chain is illustrated in the Fig. 2(c). This chain serves two main purposes: first, it facilitates MoSE’s reasoning and training; second, it provides better interpretation at inference stage. For instance, in the example shown in Fig. 2(c), the model first detects all key objects within a specific region. It then predicts the possible behaviors of these objects and finally evaluates their importance and priority. Based on this chain, the model determines which objects in the current scene are likely to influence driving decisions. This structured reasoning process offers valuable debugging cues and enhances interaction with users, thereby improving user trust in the model.
Figure 4: Pipeline for multi-modal scaling up. Building on the existing pretrained router, we further employ rule-based token allocation to increase the number of experts, thereby achieving scalability. This approach enhances the model’s capability without requiring additional data annotation or pretraining.
Figure 5: The skills definition in this paper, ensuring comprehensive coverage of QA data across multiple datasets. The marked skills represent the defined general skills.
3.1.2 Skill Data Annotation
Based on the skill definition, we randomly sample a subset of data from the target dataset and annotate them with corresponding skills. Depending on the data characteristics, we employ both rule-based and VLM-based (e.g. GPT) annotation methods (c.f. Fig. 3 Stage-2).
For relatively structured questions and general scenarios, we use a rule-based approach for annotation. These rules further divide different steps into specific skills, such as whether the question focuses on the entire image or a specific region. This method offers several advantages: it is more cost-effective, provides higher accuracy, and enables large-scale annotation at minimal expense.
For cases where rule-based methods fail to provide reliable annotations, we utilize GPT-based annotation. During this process, the input consists of the scene image, the ego vehicle’s current state, and the ground truth trajectory. Additionally, we maintain a predefined list of annotation rules, which is included as part of the prompt. These rules specify conditions such as: for incomplete samples, the general skill should be used.
During the GPT-based annotation process, we continuously refine and expand the annotation rule list both automatically and manually to improve annotation quality. As a result, higher-quality labeled data can be obtained. Further details on our annotation prompts can be found in the supplementary materials.
Although skill annotation incurs additional efforts, we show that MoSE only requires a relatively small subset of data to develop the hierarchical skill routing mechanism (c.f. Experiment Section). Furthermore, for GPT-based annotations, manual data verification is not strictly necessary, as the difficulties of both annotation and routing have already been considered during the skill definition stage.
3.1.3 Router Pretraining
After having the skill annotation data, inspired by recent work [19] we pretrain the routers in different hierarchical layers. To further enhance the efficiency, we take the mean token of all tokens from previous layer $T_{Mean}$. Taking $T_{Mean}$ as input, the training loss funciton can be expressed as Eq. (2).
$L_{Router}=-\sum_{i=1}^{n}y_{i}\log p(y_{i}|T_{Mean}).$ (2)
3.2 Expert Networks and Training
After constructing the hierarchical skill routing mechanism, we construct the MoSE layers with one shared expert and multiple skill-specialized experts. The shared expert is consistently activated to capture global information, enhancing the model’s robustness. The proposed skill centric routing mechanism ensures that the experts not only develop scenario-specific skills but also acquire distinct skills for different stages of reasoning.
Following standard practices, we adopted a sparse MoE configuration where the original FFN layers in all even-numbered Transformer layers were replaced with the proposed MoSE modules, as can be seen in Fig.  2(b). This design balances model scalability with the total number of parameters and active parameters.
We optimize the output of LLM through a generative loss in an auto-regressive manner. Given the input image and text, the model progressively generate each element $Y$. For all trainable parameter $\theta$, the loss can be expressed as:
$L_{reg}=-\sum_{i=1}^{N}\log\,p_{\theta}(Y^{[P+i]}|V,T^{[:i-1]}),$ (3)
where $Y$ is the output sequence, $K=P+N$ is the total output length. We only calculate the loss for $N$ element which is the newly generated text.
Scaling-Up on MoSE
To further enhance the model’s capabilities and scale it up without retraining the router, we propose MoSE with expert group. The overall model pipeline is illustrated the Fig. 4.
Based on the routing results from the pretrained router, we construct a group of experts, where multiple experts $n$ are assigned to $n$ sub-skills. Specifically, for current task, we adopt a modality-specific expert strategy with $n=2$ where different experts handle different modalities e.g. text or image. That’s to say, we assign two experts within each expert group: one for processing image tokens and the other for processing text tokens under the corresponding specific skill. These strategy only involved skill experts, with the shared experts remaining unchanged.
Model General Regional Suggest. Score
Closed-Source Models
Gemini-Pro [13] 25.24 51.38 27.40 34.67
GPT-4V [29] 57.50 56.26 63.30 59.02
TSLVLM+GPT4V [37] 58.70 83.41 74.26 72.12
RtoG(GPT-4o) [14] 59.00 84.37 70.80 71.39
Generalist Models
Shikra-7B [4] 12.24 22.94 10.20 15.12
LLaVA1.5-7B [25] 19.30 42.06 23.16 28.17
Qwen-VL-Chat-7B [1] 18.22 26.62 22.06 22.30
MiniCPM-V-2.5-8B [38] 41.12 57.20 48.48 48.93
LLaVA1.5-13B [25] 24.54 42.41 27.90 31.61
LLaVA-NeXT-13B [27] 29.86 53.63 31.92 38.47
InternVL-V1-5-20B [8] 38.38 61.53 41.18 47.03
Qwen-VL-Max [9] 34.60 68.17 67.40 56.72
LLaVA-one [21] 38.70 51.70 49.32 46.57
Specialist Models
CODA-8B [5] 55.04 77.68 58.14 63.62
Mini-Driver-83M [41] 21.60 62.15 45.40 43.05
Mini-Drive-137M [41] 24.60 66.34 45.44 45.46
TSLVLM-7B [37] 52.84 78.33 61.62 64.26
DriveMM-8B [16] 52.94 77.76 61.84 64.18
Drive-OV-8B [16] 53.60 77.64 58.72 63.32
Qwen2-VL_SFT 52.70 69.50 65.30 62.50
Qwen2-VL_MoE_41 55.00 71.10 66.90 64.33
Qwen2-VL_MoE_81 47.20 66.80 59.20 57.73
MoSE (Ours) 58.10 71.10 68.90 66.03
MoSE w. Exp Grp (Ours) 56.40 72.40 70.30 66.40
Table 1: Performance comparison of different models. The results in gray are based on closed-source models.The row in gray are the baseline models and the row in yellow are our proposed models. The results in bold indicate the best performance within each category. Compared to models based on open-source models and data, our approach achieves SOTA performance with significantly reduced activated model size
Figure 6: Visualization for reasoning results, where GPT Evaluation reflects the GPT model’s assessment of the generated results. The white-boxed region in the image shows the expert skills activated at different MoSE layers.
4 Experimental Results
Dataset Settings
We conduct our main experiments on CODA dataset [5] which is designed for autonomous driving reasoning task. The dataset collects driving scenarios and extract objects which may influence ego vehicle, the final driving suggestion is generated considering all key objects. Compared to other driving reasoning dataset, CODA focus on multi-modal corner cases and consider the hierarchy data structure. It require models to generate relatively long answers containing description and analyzing of multiple objects. In this paper, we follow the official training and testing settings, and use GPT-4o for evaluation. For skill annotation, following the pipeline we introduce in Sec. 3.1.1, we obtain a skill definition as in Fig. 5.
Model Settings
We experiment the MoSE based on Qwen2-VL-2B [2] which is a small-scale VLM. We adapt a total of $14$ MoSE layers. Experimental results indicate that a relatively uniform hierarchical distribution yields better performance. Therefore, in our final configuration, we allocate $[4,5,5]$ layers to perception, prediction, and planning, respectively.
The input image size is set to $[800,450]$ to balance the image quality and GPU memory. We follow the sparse setting of general MoE that only half of the layers are equipped with MoSE. During the pretraining stage and finetuning stage, we set expert number to $4$ and use $1$ shared expert. Additionally, we also experiment with general MoE method on same backbone model (denoted as Qwen2-VL_MoE in the result table), with same number of experts and number of activated experts to our MoSE. We finetune a QwenVL2-2B under a supervised finetuning settings (denoted as Qwen2-VL_SFT). For hierarchical skill routing, we annotate $2000$ skill samples using rule-based method to pretrain the routers.
Figure 7: Visualization of expert weight distribution similarity in the final layer. The expert distribution in the MoSE model is more diverse. The similarity values are log-transformed for better clarity in reflecting differences
Skill Data Size General Regional Suggest. Score
2000 58.1 68.9 71.1 66.03
3000 56.7 68.9 70.7 65.43
Table 2: Comparison on different skill-annotated data size. Additional skill data does not result in a performance gain.
Model Acc GPT Match Bleu1 Bleu2 Bleu3 Bleu4 ROUGEL CIDEr Final
MoE-StableLM 84.0 74.3 37.5 79.2 72.7 66.7 60.8 73.6 17.2 63.7
MoSE-StableLM 89.3 76.0 39.1 79.0 72.5 66.6 60.8 73.8 19.0 65.8
Table 3: QA evaluation on DriveLM-nuScences
4.1 Experiment results on CODA
Table.1 shows our results on CODA compared to other state-of-the-art models. CODA contains different subsets, general perception task lies in a comprehensive understanding of critical road key entities in driving scenarios (denoted “General”), regional perception task focus on understanding corner case objects when provided with specific bounding boxes (denoted “Regional”), and driving suggestions task aims to formulat driving advice (denoted “Suggest.”). The results are evaluated using GPT-4o, which is the same evaluation model according to CODA (denoted “Score”). The reported results contains generalist open-source and closed-source models, also specialist models which are supervised finetuned on CODA training set. As can be seen from the table, with less than 3B parameters, the proposed MoSE exceeds other SOTA methods with much less parameters. TSLVLM$+$GPT4V have the best performance but it relies on GPT for refinement.
At the bottom of the table, we compare MoSE with baseline methods, MoE, and SFT. The results indicate that incorporating MoE for scaling up the base model effectively improves performance. However, it also increases training difficulty and imposes higher demands on data.
From the results of MoE with 8 experts, the performance decline suggests that the CODA training set is insufficient to fully support the model’s training, which empirically demonstrate that general straightfoward adapting MoE encounters challenges and often results in suboptimal performance. In contrast, the proposed MoSE outperforms MoE. Further, the MoSE model with expert groups, despite having the same training data as MoE-8-1, does not suffer from performance degradation. Instead, it achieves a slight improvement, which we attribute to effective data utilization bringing by hierarchical skill routing.
Beyond the performance improvements demonstrated by the evaluation results on driving suggesting task, One thing need to be mentioned is that, the official setting of CODA uses general result and object result of context to generate final driving suggestion. In our settings, we perform single round QA without multi-turn conversation and long context.
In Fig.7 we visualize the similarity of each expert in MoE layer. During the visualization, we visualize the last MoE layer, and analyze the weights of one of the MLP layers. In a MoE framework, greater differentiation among experts enhances model performance by promoting specialization and efficient resource allocation. When experts are highly distinct, each one is able to focus on a specific subset of tasks or input distributions, leading to more refined representations and better decision-making.This differentiation helps mitigate redundancy, ensuring that experts contribute unique knowledge rather than overlapping in functionality. Additionally, a more diverse set of experts reduces competition for the same data, allowing the router to make more discriminative assignments, improving both training efficiency and generalization. As show in the Fig.7, with the hierarchical skill routing strategy, our experts have more diverse weights. For visualization of other layers, please refer to supplementary.
Figure 8: Performance variations of MoE and MoSE after training on different data scales. MoSE demonstrates better performance scaling than MoE as data size increases, highlighting its potential for larger and more complex datasets.
4.2 Further Analysis
Training data size
In Fig.8,we discussed model performance with different training data-size. During the training, we randomly sample half and quarter of general, object and driving suggestion QAs from CODA and train MoE, MoSE under the same setting with same dataset, and use the official testing set for evaluation. As the data scale increases, we observe that the performance of the MoE model peaks when trained with half data, with no further improvements as the data size continues to grow. We hypothesize that this plateau may be attributed to the way experts are allocated.
In contrast, the model trained with hierarchical skill routing exhibits a steady improvement in performance. This has been validated in both versions of the proposed MoSE. This suggests that as the dataset size increases, MoSE has the potential to achieve even better training results.
Data Size for Skill Annotation
Given that the proposed MoSE, compared to the original MoE, relies on a small annotated data subset, we may wonder whether this additional requirement could limit the model’s applicability to other datasets or tasks, or whether it would hinder its scalability.
To address this, we design experiments to demonstrate that MoSE does not depend on large-scale additional data for pretraining but instead requires only a tiny subset. In the experiments shown in Table.2, we pretrain the model’s router using $2000$ and $3000$ skill-labeled samples, respectively. The results indicate that further scaling up the skill dataset does not lead to additional performance gains, thereby confirming that MoSE is not constrained by the need for extensive additional annotations.
4.3 Extension on other dataset
In this section, we further conduct experiment on DriveLM dataset with different backbone and skill settings. Compared to CODA, DriveLM dataset is based on nuScense dataset containing more common scenarios. Meanwhile, DriveLM considers more about driving planning and trajectory estimation, and has more diverse questions.
In the experiments, we follow a different skill settings and build a MoSE based on StableLM-1.6B[2]. For each driving scenario, we set one skill based on the scene itself, and build skill dataset using GPT-4o. In DriveLM, six-view images are used as inputs. Considering the token length limitation of the backbone model, we pair the images two by two before inputting them to reduce the number of tokens. During the experiments, we randomly select one-sixth of the original training set as the test set, while the remaining data is used for training.
Experiment results are shown in Table. 3. As can be seen in the results, the proposed MoSE has better performance with different settings in the DriveLM dataset, which once again prove the effectiveness of the proposed MoSE, and show that it can be easily extended to other tasks. (See Supplementary for more results).
5 Conclusion
In this paper, we proposed a task oriented Mixture-of-Experts method named Mixture-of-Skill-Experts, which mimic human learning and reasoning process. With the proposed hierarchical skill routing strategy, the model learns from a dataset step-by-step, skill-by-skill. Empirically, MoSE demonstrates that a thoughtfully designed, skill-oriented routing mechanism can achieve superior performance in autonomous driving tasks. It effectively balances model complexity, training efficiency, and data requirements, offering a promising direction for future developments in the field. Besides, a scaling-up strategy is proposed to further boost the model’s capacity without additional data or pretraining. Further experiments show that MoSE’s hierarchical skill routing not only enhances performance with increasing data sizes but also maintains scalability by necessitating only a minimal amount of additional layers within the expert groups.
Limitation and Future work
In this work, our MoSE model is trained on a relatively small subset of available data, which means the resulting performance may still have potential to be improved. Currently, our model is designed primarily for reasoning-related tasks without direct trajectory estimation. Therefore, further experimentation is necessary to investigate whether MoSE can also benefit trajectory estimation. Additionally, integrating trajectory estimation with closed-loop evaluations, such as in simulation environments like Carla, could further enhance the practical impact of MoSE. Similar to general MoE approaches, MoSE is easily adaptable to various pre-trained models for fine-tuning on diverse downstream tasks. Such directions could be further pursued in the future.